# Can LLMs Generate Diverse Molecules? Towards Alignment with Structural Diversity

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have demonstrated impressive performance in generating molecular structures as drug candidates, which offers significant potential to accelerate drug discovery. However, the current LLMs overlook a critical requirement for drug discovery: proposing a diverse set of molecules. This diversity is essential for improving the chances of finding a viable drug, as it provides alternative molecules that may succeed where others fail in wet-lab or clinical validations. Despite such a need for diversity, the LLMs often output structurally similar molecules from a given prompt. While decoding schemes like beam search may enhance textual diversity, this often does not align with molecular structural diversity. In response, we propose a new method for fine-tuning molecular generative LLMs to autoregressively generate a set of structurally diverse molecules, where each molecule is generated by conditioning on the previously generated molecules. Our approach consists of two stages: (1) supervised fine-tuning to adapt LLMs to autoregressively generate molecules in a sequence and (2) reinforcement learning to maximize structural diversity within the generated molecules. Our experiments show that (1) our fine-tuning approach enables the LLMs to better discover diverse molecules compared to existing decoding schemes and (2) our fine-tuned model outperforms other representative LLMs in generating diverse molecules, including the ones fine-tuned on chemical domains.
@@ 1 Introduction
Figure 1: LLMs generate a molecule with SMILES representation.
Recent advances in large language models (LLMs) have demonstrated remarkable potential to accelerate scientific discovery by leveraging their language processing capabilities. This progress has been particularly impactful for candidate design problems such as drug discovery (Pei et al., 2024), protein sequence design (Zhuo et al., 2024), and material design (Gruver et al., 2024). In particular, with expansive biomolecular datasets and molecular string representations, e.g., SMILES (Weininger, 1988) or SELFIES (Krenn et al., 2020), LLMs have demonstrated impressive abilities to generate molecules from textual descriptions, e.g., molecular properties (Edwards et al., 2022; Ye et al., 2023; Pei et al., 2024), as illustrated in Figure 1.
However, current LLM-based molecular generation approaches (Edwards et al., 2022; Ye et al., 2023; Pei et al., 2024) often overlook a critical requirement for drug discovery: proposing a diverse set of molecules. In computer-aided drug discovery pipelines, identifying a single molecule with a desired property does not guarantee success in real-world pipelines that require additional cell-based studies and clinical trials (Vamathevan et al., 2019). Therefore, drug discovery requires a collection of structurally diverse molecules, as illustrated in Figure 2. The generation of structurally diverse molecules increases the chances of finding a viable drug candidate (Xie et al., 2023), as different molecules may succeed where others fail. This diversity is essential to enhance the robustness and success of the drug discovery process (Krantz, 1998; Hong et al., 2020; Sadybekov & Katritch, 2023).
Figure 2: The compute-aided drug design. A collection of structurally diverse molecules is required to increase the chance of identifying viable drug candidates in the real world.
Existing LLMs have limitations in generating diverse molecules. To obtain diverse molecules, one may consider querying the recently developed generalist LLMs, e.g., Llama (Touvron et al., 2023) or ChatGPT (OpenAI, 2023). However, our empirical observation in Figure 3(a) reveals that even the most recent models produce structurally identical or highly similar molecules from the given prompt. [^1] ChatGPT-4o (OpenAI, 2023) generates different SMILES strings that map to an identical molecule. This observation aligns with previous observations that have shown LLMs may fail to generate diverse outputs (Kirk et al., 2024) for general text-based domains.
@@
Our approach. We repurpose existing molecular generative LLMs to autoregressively generate a diverse set of molecules from a single prompt. By enabling the LLMs to generate a new molecule conditioned on previously generated molecules, we expect the LLMs to learn to enhance the structural diversity between the generated molecules. To this end, we propose a two-stage approach to fine-tune LLMs: (a) a supervised fine-tuning stage to repurpose LLMs to autoregressively generate a sequence of multiple molecules and (b) a reinforcement learning stage to maximize the molecular structural diversity. Note that both stages do not require external diverse molecular datasets and are purely based on self-improvement procedures, where the LLMs train on the samples generated by themselves.
In the supervised training stage, we train the LLMs to autoregressively generate a set of molecules in a single sequence. The training dataset, i.e., a set of molecules, is collected from the LLMs themselves through iterative sampling, and then filtered to enhance the quality, e.g., removing invalid molecules. However, this stage does not necessarily incorporate molecular diversity, as the training may not involve sufficiently distinct molecules (e.g., limitations in Figure 3(b)). To tackle this, we subsequently apply reinforcement learning with exploration towards discovering diverse molecules.
@@
We compare our method with the decoding schemes for diversified generation (Vijayakumar et al., 2016; Su et al., 2022) and other representative LLMs, including chemical-task specialists (Edwards et al., 2022; Christofidellis et al., 2023; Pei et al., 2023; 2024), fine-tuned generalists on chemical domains (Fang et al., 2024; Yu et al., 2024), and the ChatGPT series (OpenAI, 2023; 2024). We observe that (1) our fine-tuning approach enables LLMs to better discover diverse molecules compared to existing decoding schemes and (2) our fine-tuned LLM outperforms other recent LLMs.
@@
  • Experimentally, our method outperforms the baselines in generating diverse molecular structures.
@@ 2 Related work
Large language models (LLMs) for molecular generation. Recent advancements in LLMs have shown increasing promise in scientific applications, especially for molecular generation (Edwards et al., 2022; Pei et al., 2023; Fang et al., 2024; Pei et al., 2024). First, Edwards et al. (2022) proposed MolT5 which translates between SMILES (Weininger, 1988) and molecular text descriptions. Text+Chem T5 (Christofidellis et al., 2023) is a model pre-trained on both the chemical and natural language domains. Next, BioT5 (Pei et al., 2023) considers T5 models pre-trained on datasets including bio-text, protein sequences, and molecules. Additionally, Ye et al. (2023), Fang et al. (2024), and Yu et al. (2024) fine-tuned generalist LLMs, e.g., Llama (Touvron et al., 2023), through biological instructions, molecular modifications, and large-scale molecular datasets, respectively.
Decoding schemes for generating diverse output sequences. To generate diverse and high-quality solution candidates from LLMs, existing literature on LLMs has studied improving decoding schemes. To acquire multiple distinct sequences with high likelihoods, one can consider employing beam search, which jointly decodes multiple distinct outputs (Och, 2003). To enhance diversity between sequences, Vijayakumar et al. (2016; 2018) incorporated token-wise differences between generated sequences in the beam search. Furthermore, Su et al. (2022) considered the contrast between the candidate sequences. In addition, Holtzman et al. (2020) proposed nucleus sampling, which enhances random sampling by balancing the quality and the diversity.
Reinforcement learning (RL) for fine-tuning LLMs. RL has been effectively applied to fine-tune LLMs, aligning them with desired behaviors expressed through reward signals. One notable example is RL from human feedback to align LLMs with human preference (Ouyang et al., 2022). In addition, there has been a surge in research on devising RL for LLMs as well, such as addressing multi-turn settings (Shani et al., 2024) and incorporating multiple fine-grained reward signals (Wu et al., 2023b). For molecular generation, Ghugare et al. (2024) proposed RL-based fine-tuning to generate a molecule satisfying target properties. However, to the best of our knowledge, there exist no prior RL-based approaches that aim to increase the diversity of LLM-generated outputs.
(a) Supervised fine-tuning: enables generating a sequence of multiple molecules.
(b) Reinforcement learning: maximizes the reward of generated molecules.
(c) Reward computation: captures the structural diversity.
Figure 4: Illustration of proposed fine-tuning approaches. We consider two stages for fine-tuning LLMs: a supervised fine-tuning Figure 4(a) and a reinforcement learning Figure 4(b). The prompts are simplified for explanatory purposes, and the actual prompts are provided in Appendix B.
@@ 3 Method
In this section, we present our method for fine-tuning LLMs to generate diverse molecules. Specifically, we consider applying fine-tuning to existing molecular generative LLMs that produce molecular representations such as SMILES or SELFIES. Importantly, our approach is versatile and can be broadly applied to other domains, e.g., protein sequence (Zhuo et al., 2024) or computer-aided design (Wu et al., 2023a). Furthermore, it leverages self-improvement techniques and does not require additional datasets containing diverse molecules.
Overview. Our goal is to generate a sequence of structurally diverse molecules from a given prompt by producing them in a single concatenated output. To achieve this, we fine-tune the LLMs in two stages: (a) a supervised fine-tuning phase that repurposes the LLMs to generate a sequence of molecules rather than a single one, and (b) a reinforcement learning phase aimed at further enhancing the structural diversity among the generated molecules.
Task details. In detail, we consider generating molecules from a prompt $p_{\text{desc}}$, where the prompt describes a molecular property that the generated molecules should possess. Here, we aim to generate diverse molecules that satisfy the given description $p_{\text{desc}}$. We let $\mathcal{P}$ denote the prompts used for training, which does not intersect with the ones used for evaluation.
@@
3:     Sample prompt $p_{\text{desc}}\sim\mathcal{P}$
4:     Sample $\{m_{i}\}_{i=1}^{T}$ from $\pi_{\text{pre}}(m\;|\;p_{\text{desc}})$
@@ 3.1 Supervised fine-tuning
Dataset collection. The dataset collection and supervised training processes are conducted with a set of training prompts $\mathcal{P}$. Initially, the pre-trained LLM $\pi_{\text{pre}}$ produces a set of molecules by iterative sampling molecules for a given prompt $p_{\text{desc}}\in\mathcal{P}$ as follows:
@@
where $m_{i}$ denotes the string representation of the molecule. In practice, we employ beam search to collect the set of molecules $\{m_{i}\}^{T}_{i=1}$. Then, we filter out the invalid string representations, duplicate molecules, and molecules that do not satisfy the given prompt $p_{\text{desc}}$. This results in reducing the set of molecules from $\{m_{i}\}^{T}_{i=1}$ to $\{m_{i}\}^{K}_{i=1}$. The details are described in Appendix A.
@@ 3.2 Reinforcement learning
5:     Sample $m_{k}\sim\pi_{\text{RL}}(m_{k}\mid\mathcal{M}_{1:k-1},p_{\text{desc+div}})$
@@
We apply reinforcement learning to maximize the diversity of the generated molecules within a sequence. However, when applied to a sequence of molecules $\mathcal{M}_{1:K}$, conventional sequence-wise reinforcement learning (Ouyang et al., 2022) suffers from the credit assignment problem (Zhou et al., 2024): the challenge in identifying and promoting the generation of molecules responsible for increasing diversity, among a set of molecules $\{m_{i}\}^{K}_{i=1}$. To circumvent this, we introduce a molecule-wise reinforcement learning. [^2] We compare both approaches in Table 4 of Section 4.4.
Specifically, we consider reinforcement learning on a sequence of molecules $\mathcal{M}_{1:K}$ as learning in $K$ individual stages. Each stage corresponds to generating a molecule $m_{k}$ conditioned on a sequence of previously generated molecules $\mathcal{M}_{1:k-1}$. Then, the LLM $\pi_{\text{RL}}$ is trained to maximize the return of each stage, which is defined by the reward of the generated molecule $m_{k}$. The reward is defined as the diversity between the previously generated molecules $\{m_{i}\}_{i=1}^{k-1}$ and the new molecule $m_{k}$. We also incorporate an auxiliary reward to ensure that the molecule $m_{k}$ satisfies the given description $p_{\text{desc}}$. We illustrate our approach in Figure 4(b) and provide the detailed algorithm in Algorithm 2.
Reward. The reward evaluates the generated molecule $m_{k}$ with a diversity reward $r_{\text{div}}(m_{k},\{m_{i}\}^{k-1}_{i=1})$ and a description-matching reward $r_{\text{match}}(m_{k},p_{\text{desc}})$, as follows:
@@
where the diversity reward $r_{\text{div}}$ evaluates structural differences between the molecule $m_{k}$ and the previously generated molecules $\{m_{i}\}^{k-1}_{i=1}$ by assessing their true molecular structures. Note that $r_{\text{div}}(m_{1})$ is zero. The description-matching reward $r_{\text{match}}$ evaluates whether the molecule $m_{k}$ satisfies the given description $p_{\text{desc}}$. In a practical implementation, the final action that completes the molecule $m_{k}$ yields the reward $r(m_{k})$. The detailed reward implementation is described in Appendix A.
Policy optimization. We optimize the LLM $\pi_{\text{RL}}$ to maximize the reward using proximal policy optimization (PPO; Schulman et al., 2017; Ouyang et al., 2022). Note that $\pi_{\text{RL}}$ is initialized with $\pi_{\text{SFT}}$. In addition, we combine per-token KL penalty from the supervised fine-tuned model at each token following prior studies (Ouyang et al., 2022).

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have demonstrated impressive performance in molecular generation, which offers potential to accelerate drug discovery. However, the current LLMs overlook a critical requirement for drug discovery: proposing a diverse set of molecules. This diversity is essential for improving the chances of finding a viable drug, as it provides alternative molecules that may succeed where others fail in real-world validations. Nevertheless, the LLMs often output structurally similar molecules. While decoding schemes like diverse beam search may enhance textual diversity, this often does not align with molecular structural diversity. In response, we propose a new method for fine-tuning molecular generative LLMs to autoregressively generate a set of structurally diverse molecules, where each molecule is generated by conditioning on the previously generated molecules. Our approach consists of two stages: (1) supervised fine-tuning to adapt LLMs to autoregressively generate molecules in a sequence and (2) reinforcement learning to maximize structural diversity within the generated molecules. Our experiments show that the proposed approach enables LLMs to generate diverse molecules better than existing approaches for diverse sequence generation.
Can LLMs Generate Diverse Molecules?
Towards Alignment with Structural Diversity
Hyosoon Jang1, Yunhui Jang2, Jaehyung Kim3, Sungsoo Ahn2 1POSTECH 2KAIST 3Yonsei University hsjang1205@postech.ac.kr, jaehyungk@yonsei.ac.kr, {yunhuijang,sungsoo.ahn}@kaist.ac.kr
@@ 1 Introduction
Recent advances in large language models (LLMs) have demonstrated the potential to accelerate scientific discovery by leveraging their language processing capabilities. This progress has been particularly impactful for candidate design problems such as drug discovery (Pei et al., 2024), protein design (Zhuo et al., 2024), and material design (Gruver et al., 2024). In particular, with biomolecular datasets and molecular string representations, e.g., SMILES (Weininger, 1988) or SELFIES (Krenn et al., 2020), LLMs have demonstrated impressive abilities to generate molecules from textual descriptions, e.g., molecular properties (Edwards et al., 2022; Ye et al., 2023; Pei et al., 2024).
However, current LLM-based molecular generation approaches (Edwards et al., 2022; Ye et al., 2023; Pei et al., 2024) often overlook a critical requirement for drug discovery: proposing a diverse set of molecules. In computer-aided drug discovery, identifying a single molecule with a desired property does not guarantee success in real-world pipelines that require additional cell-based studies and clinical trials (Vamathevan et al., 2019). Therefore, drug discovery requires a set of structurally diverse molecules. The generation of structurally diverse molecules increases the chances of finding a viable drug candidate (Xie et al., 2023), as different molecules may succeed where others fail. This diversity is essential to enhance the robustness and success of the drug discovery (Krantz, 1998; Hong et al., 2020; Sadybekov and Katritch, 2023).
@@
Existing LLMs have limitations in generating diverse molecules. To obtain diverse molecules, one may query the recent generalist LLMs, e.g., Llama (Touvron et al., 2023) or ChatGPT (OpenAI, 2023). However, our empirical observation in Figure 1(a) reveals that even the recent models produce structurally identical or highly similar molecules from the given prompt. [^1] ChatGPT-4o (OpenAI, 2023) generates different SMILES strings that map to an identical molecule. This observation aligns with previous observations that have shown LLMs may fail to generate diverse outputs (Kirk et al., 2024) for general text-based domains.
Our approach. We repurpose existing molecular generative LLMs to autoregressively generate a diverse set of molecules from a single prompt. By enabling the LLMs to generate a new molecule conditioned on previously generated molecules, we expect the LLMs to learn to enhance the structural diversity between the generated molecules. To this end, we propose a two-stage approach to fine-tune LLMs: (a) a supervised fine-tuning stage to repurpose LLMs to autoregressively generate a sequence of multiple molecules and (b) a reinforcement learning stage to maximize the molecular structural diversity between the generated molecules.
In the supervised training stage, we train LLMs to autoregressively generate a set of molecules in a sequence. Note that the training data, i.e., a set of molecules, can be collected from LLMs themselves through iterative sampling, and then filtered to enhance the quality, e.g., removing invalid molecules. However, this stage does not necessarily incorporate molecular diversity, as the training may not involve sufficiently distinct molecules (e.g., limitations in Figure 1(b)). To tackle this, we subsequently apply reinforcement learning with exploration towards discovering diverse molecules.
@@
We compare our method with the decoding schemes for diversified generation (Vijayakumar et al., 2016; Su et al., 2022) and other representative LLMs, including chemical-task specialists (Edwards et al., 2022; Christofidellis et al., 2023; Pei et al., 2023, 2024), fine-tuned generalists on chemical domains (Fang et al., 2024; Yu et al., 2024), and the ChatGPT series (OpenAI, 2023, 2024). We observe that (1) our fine-tuning approach enables LLMs to better discover diverse molecules compared to existing decoding schemes and (2) our fine-tuned LLM outperforms other LLMs.
@@
  • Experimentally, our method outperforms the baselines in generating diverse molecules.
@@ 2 Related Work
Large language models (LLMs) for molecular generation. Recent advancements in LLMs have shown increasing promise in scientific applications, especially for molecular generation (Edwards et al., 2022; Pei et al., 2023; Fang et al., 2024; Pei et al., 2024). First, Edwards et al. (2022) proposed MolT5, a molecular generative LLM that translates between SMILES (Weininger, 1988) and molecular text descriptions. Next, Text+Chem T5 (Christofidellis et al., 2023) and BioT5 (Pei et al., 2023) considered pre-training molecular generative LLMs on datasets that incorporate extensive chemical knowledge, e.g., scientific articles. Additionally, Ye et al. (2023), Fang et al. (2024), (Chen et al., 2023), and Yu et al. (2024) fine-tuned generalist LLMs, e.g., Llama (Touvron et al., 2023), through biological instructions, molecular modifications, and large-scale molecular datasets, respectively.
Decoding schemes for generating diverse output sequences. To generate diverse solution candidates from LLMs, existing literature on LLMs has studied improving decoding schemes. To acquire multiple distinct sequences with high likelihoods, one can consider employing beam search, which jointly decodes multiple distinct outputs (Och, 2003). To enhance diversity between sequences, Vijayakumar et al. (2016, 2018) incorporated token-wise differences between generated sequences in the beam search. Furthermore, Su et al. (2022) considered the contrast between the candidate sequences. In addition, Holtzman et al. (2020) proposed nucleus sampling, which enhances random sampling by balancing the quality and the diversity.
Reinforcement learning (RL) for fine-tuning LLMs. RL has been effectively applied to fine-tune LLMs, aligning them with desired behaviors expressed through reward signals. One notable example is RL from human feedback to align LLMs with human preference (Ouyang et al., 2022). In addition, there has been a surge in research on devising RL for LLMs as well, such as addressing multi-turn settings (Shani et al., 2024) and incorporating multiple fine-grained reward signals (Wu et al., 2023b). For molecular generation, Ghugare et al. (2024) proposed RL-based fine-tuning to generate a molecule satisfying target properties. However, to the best of our knowledge, there exist no prior RL-based approaches that aim to increase the diversity of LLM-generated outputs. [^2] More related works, e.g., RL for diverse molecular generation without using LLMs, are described in Appendix A.
(a) Supervised fine-tuning (Algorithm 1): enables generating a sequence of multiple molecules.
(b) Reinforcement learning (Algorithm 2): maximizes the diversity between new and previously generated molecules.
Figure 2: Illustration of proposed fine-tuning approaches. We consider two stages for fine-tuning LLMs: a supervised fine-tuning Figure 2(a) and a reinforcement learning Figure 2(b). The prompts are simplified for explanatory purposes, and the actual prompts are provided in Appendix D.
@@ 3 Method
In this section, we present our method for fine-tuning LLMs to generate diverse molecules. Specifically, we consider fine-tuning existing molecular generative LLMs that produce molecular representations such as SMILES or SELFIES. Importantly, our approach is versatile and can be applied to other domains, e.g., protein sequence (Zhuo et al., 2024) or computer-aided design (Wu et al., 2023a).
Overview. Our goal is to generate a sequence of structurally diverse molecules from a given prompt by producing them in a single concatenated output. To this end, we fine-tune the LLMs in two stages: (a) a supervised fine-tuning phase that repurposes the LLMs to generate a sequence of molecules rather than a single one, and (b) a reinforcement learning phase aimed at further enhancing the structural diversity among the generated molecules.
Task details. In detail, we consider generating molecules from a prompt $p_{\text{desc}}$, where the prompt describes a molecular property that the generated molecules should possess. In this setting, we aim to generate diverse molecules that satisfy the given description $p_{\text{desc}}$, where the diversity is evaluated using similarity measures between the structural features of the molecules, e.g., the presence of specific atoms, or substructures (Bajusz et al., 2015). We let $\mathcal{P}$ denote the prompts used for training.
3.1 Supervised Fine-tuning
@@
3:     Get prompt $p_{\text{desc}}\sim\mathcal{P}$
4:     Get $\{m_{i}\}_{i=1}^{T}$ from $\pi_{\text{pre}}(m\;|\;p_{\text{desc}})$
@@
Dataset collection. The supervised training process is conducted with a set of training prompts $\mathcal{P}$. Initially, the pre-trained LLM $\pi_{\text{pre}}$ produces a set of molecules by iterative sampling molecules for a given prompt $p_{\text{desc}}\in\mathcal{P}$ as follows:
@@
where $m_{i}$ denotes the string representation of the molecule. In practice, we employ beam search to collect the set of molecules $\{m_{i}\}^{T}_{i=1}$. Then, we filter out the invalid string representations, duplicate molecules, and molecules that do not satisfy the given prompt $p_{\text{desc}}$. This results in reducing the set of molecules from $\{m_{i}\}^{T}_{i=1}$ to $\{m_{i}\}^{K}_{i=1}$. The details are described in Appendix B.
@@ 3.2 Reinforcement Learning
5:     Get $m_{k}\sim\pi_{\text{RL}}(m_{k}\mid\mathcal{M}_{1:k-1},p_{\text{desc+div}})$
@@
Table 1: Comparison with existing decoding schemes. NCircles represents both quality and diversity-related metric. Accepted & unique represent quality-related metrics. IntDiv. represents an average of pair-wise diversities. Our method (Div-SFT+RL) generates more diverse and high-quality molecules compared to the baselines. Notably, our method makes a larger gap over the baselines on NCircles related to capturing both quality and diversity.
Dataset Method NCirclesh=0.85 NCirclesh=0.75 Accepted & Unique IntDiv.
Random 1.079 0.948 1.970 0.109
Nucleus 1.006 0.918 1.623 0.090
BS 4.562 2.603 33.406 0.176
L+M-24 Diverse BS 2.395 1.743 9.915 0.234
Contrastive BS 4.521 2.594 33.568 0.176
Div-SFT 5.198 3.205 20.711 0.250
Div-SFT+RL 10.51 6.278 31.060 0.287
Random 1.539 1.383 1.998 0.045
Nucleus 0.897 0.876 0.972 0.014
BS 5.576 4.171 11.734 0.194
ChEBI-20 Diverse BS 4.413 3.637 5.905 0.140
Contrastive BS 7.955 5.883 15.215 0.233
Div-SFT 4.869 3.601 8.178 0.141
Div-SFT+RL 11.301 8.996 16.271 0.246
We apply reinforcement learning to maximize the diversity of the generated molecules within a sequence. However, when applied to a sequence of molecules $\mathcal{M}_{1:K}$, conventional sequence-wise reinforcement learning (Ouyang et al., 2022) suffers from the credit assignment problem (Zhou et al., 2024): the challenge in identifying and promoting the generation of molecules responsible for increasing diversity. To circumvent this, we introduce a molecule-wise reinforcement learning. [^3] We make comparisons between the sequential- and the molecule-wise approaches in Table 5 of Section 4.4.
Specifically, we consider reinforcement learning on a sequence of molecules $\mathcal{M}_{1:K}$ as learning in $K$ individual stages. Each stage corresponds to generating a molecule $m_{k}$ conditioned on a sequence of previously generated molecules $\mathcal{M}_{1:k-1}$. Then, the LLM $\pi_{\text{RL}}$ is trained to maximize the return of each stage, defined by the reward of the generated molecule. Here, the reward is the diversity between the previously generated molecules $\{m_{i}\}_{i=1}^{k-1}$ and the new molecule $m_{k}$. We also incorporate an auxiliary reward to ensure that the generated molecule satisfies the description $p_{\text{desc}}$. We present our approach in Figures 2(b) and 2.
Reward. The reward evaluates the molecule $m_{k}$ with a diversity reward $r_{\text{div}}(m_{k},\{m_{i}\}^{k-1}_{i=1})$ and a description-matching reward $r_{\text{match}}(m_{k},p_{\text{desc}})$:
@@
where the diversity reward $r_{\text{div}}$ evaluates structural differences between the molecule $m_{k}$ and the previously generated molecules $\{m_{i}\}^{k-1}_{i=1}$ by assessing their true molecular structures. Note that $r_{\text{div}}(m_{1})$ is zero. The description-matching reward $r_{\text{match}}$ evaluates whether the molecule $m_{k}$ satisfies the description $p_{\text{desc}}$. In practice, the final action completing the molecule yields the reward. The implementation is described in Section 4.
Policy optimization. We optimize the LLM $\pi_{\text{RL}}$ to maximize the reward using proximal policy optimization (PPO; Schulman et al., 2017). Here, $\pi_{\text{RL}}$ is initialized with $\pi_{\text{SFT}}$. We also combine per-token KL penalty from the supervised fine-tuned model following prior studies (Ouyang et al., 2022).
Table 2: Visualization of the generated molecules. We generate ten molecules from: “The molecule is a egfr inhibitor and modulate, belonging to the cancer treatment class of molecules, and is treatment of disorder”. The blue line indicates the molecule that follows the given description, as evaluated using the external tool (Trott and Olson, 2010). BS generates structurally similar molecules. The details are described in Appendix E.
BS Ours
[Uncaptioned image] [Uncaptioned image]